OpenStudy (anonymous):

5^7x-2=4^x solve the equation

OpenStudy (anonymous):

\[\LARGE 5^{7x-2}=4^x\] ?

OpenStudy (anonymous):

5^(7x-2)=4^(x) ln5^(7x-2)=ln 4^(x) 7xln5 -2ln 5- xln4 = 0 x= 2ln 5 / 7ln 5- ln4 x = aprox. 3.2189

OpenStudy (anonymous):

i have another on its 2^4x+2=8^x